# Learning to Drive Anywhere with Model-Based Reannotation

### v1

@@ Abstract
Developing broadly generalizable visual navigation policies for robots is a significant challenge, primarily constrained by the availability of large-scale, diverse training data. While curated datasets collected by researchers offer high quality, their limited size restricts policy generalization. To overcome this, we explore leveraging abundant, passively collected data sources, including large volumes of crowd-sourced teleoperation data and unlabeled YouTube videos, despite their potential for lower quality or missing action labels. We propose Model-Based ReAnnotation (MBRA), a framework that utilizes a learned short-horizon, model-based expert model to relabel or generate high-quality actions for these passive datasets. This relabeled data is then distilled into LogoNav, a long-horizon navigation policy conditioned on visual goals or GPS waypoints. We demonstrate that LogoNav, trained using MBRA-processed data, achieves state-of-the-art performance, enabling robust navigation over distances exceeding 300 meters in previously unseen indoor and outdoor environments. Our extensive real-world evaluations, conducted across a fleet of robots (including quadrupeds) in six cities on three continents, validate the policy’s ability to generalize and navigate effectively even amidst pedestrians in crowded settings. We open-source our models and codes and provide supplementary videos on our project page  [^1] https://model-base-reannotation.github.io/
Keywords: vision-based navigation, annotation, model-based learning
Figure 1: We train a highly generalizable navigation policy that can control robots in a variety of conditions and be deployed zero-shot in new environments across the world. Our proposed method, Model-Based ReAnnotation, enables imitation learning from noisy, passive data, such as low-quality crowd-sourced demonstrations or even videos from the web.
1 Introduction
Machine learning has demonstrated remarkable success across a range of tasks, including natural language processing [1, 2] and computer vision [3, 4, 5]. A key factor driving these advancements is the availability of large and diverse training datasets. In robotics, lack of data is a major bottleneck: intentional, centralized data-collection efforts are costly, requiring real-world robots and human operators, while Internet-scraped data is rarely directly applicable to the robotics domain [6, 7].
In this paper, we study the problem of developing an end-to-end robot navigation policy capable of generalizing to a wide range of outdoor and indoor environments and navigating to distant goals hundreds of meters away. Training such an end-to-end policy requires large amounts of diverse data to grant broad coverage over possible environments. Previous navigation works [8] have relied on centrally collected datasets generated by robotics researchers. While these datasets tend to be high quality, the sum total of these datasets is on the order of dozens of hours [9], limiting the breadth of generalization that can be achieved from this high-quality data alone.
@@ 1 Introduction
2 Related work
Vision-based robot navigation has been widely explored to navigate toward goal positions given visual observations from a monocular camera. [10, 11, 12] train short-horizon policies to generate actions with access to a single goal observation. These short-horizon policies often utilize topological memory to extend the range of navigation [13]. Some works[14] use exploration with a topological memory to seek out a distant image goal, while others[15, 16] use a GPS signal for localization and navigate toward a goal provided as a 2D position in cartesian coordinates. Goal images and poses require prior access to the target environment and knowledge of the environment’s geometry. Various learning methodologies such as imitation learning (IL) [10, 8, 9], reinforcement learning (RL) [16, 17, 18], and model-based learning (MBL) [12, 19] have been explored for training goal-conditioned vision-based policies on publicly available robot datasets.
These methods require a sequence of image observations and corresponding actions parsed from accurate wheel odometry [12, 20], GPS [14], and other reliable sensors. These datasets are collected via intentional, centralized teleoperation efforts with the downstream goal of training a navigation policy and, therefore, contain goal-directed trajectories. Collecting data of this sort at a global scale would require a massive unified effort that would be costly and time-consuming.
Robot learning with passive data. Visual SLAM [21] and inverse dynamics models [22] can be used to estimate trajectories for first-person videos, allowing us to train policies that use these trajectories as approximations of robot actions from action-free and non-robot data. While visual SLAM and its successors [23, 24, 25, 26] offer good local trajectory estimation, its accuracy relies on having consistent, good visual features in the image view.
Robotic foundation models (RFMs) [9, 27] trained with IL on curated data can address an embodiment gap issue and augment passive data sources with consistent robotic actions [28]. However, current RFMs still lack coverage of diverse environments and cannot leverage passive data with noisy action labels during training. To address these issues, we train an expert relabeling MBRA model with MBL to better approximate reasonable robot actions. Since MBL is robust to noisy action labels during training, we can train the MBRA model with passive data sources and use it to reannotate large amounts of passive data with a smaller embodiment gap. We can then train the LogoNav, which can successfully perform a diverse set of long-distance navigation tasks and demonstrates an explicit advantage against the baseline policies.
@@ 2 Related work
Figure 2: Overview of MBRA. We propose a two-step process: In the first stage, we train a short-horizon reannotation policy with a robust MBL approach on the noisy dataset, which can be used for short-horizon image-conditioned navigation and which we leverage to relabel the noisy dataset with improved action labels. In step 2, we train a long-horizon navigation policy with the generated action labels.
@@
3 Learning Short-Horizon Relabeling Policies with Model-based Learning
@@
In this paper, we focus on a learning long-horizon navigation policy from a highly diverse but suboptimal dataset $\mathcal{D}_{n}$. In particular, we wish to learn high-quality navigation from a crowd-sourced dataset; this requires us to train a relabeler that can predict actions that are better than those found in the original dataset. We assume access to a smaller clean dataset $\mathcal{D}^{*}$ that contains high-quality behavior, such that $\lvert\mathcal{D}^{*}\rvert<\!\!<\lvert\mathcal{D}_{n}\rvert$. While observations in $\mathcal{D}_{n}$ might represent high state coverage, the actions are low-quality: both because of inaccuracy due to state estimation errors and heterogeneity of uncurated human operators with varying skill levels. Our proposed approach is a two-step process (Fig. 2): Step 1, we learn a model-based re-annotation model using MBL to learn from noisy data, and Step 2, we train a long-range navigation policy to imitate actions relabeled by the first step.
3.1 Learning a Short-Horizon Relabeling Model, MBRA
We train a model $\{a^{s}_{i}\}_{i=0\ldots N-1}=\pi^{s}(O_{c};O_{g})$, which we call the MBRA model, to infer the optimal actions occurring between the current observation $O_{c}$ and the goal observation $O_{g}$ via a model predictive control (MPC)-inspired learning approach. Our approach directly optimizes an objective function rather than imitating the dataset actions. We train $\pi^{s}$ on the joint dataset $\mathcal{D}_{n}\cup\mathcal{D}^{*}$.
We use the following model-based objective for learning $\pi^{s}$, following prior work Hirose et al. [19]:
@@ 3.1 Learning a Short-Horizon Relabeling Model, MBRA
where $s^{ref}$ is the target state and $\{s^{s}_{i}\}_{i=0\ldots N-1}$ are the estimated states at each step. $s^{s}_{i}$ is defined by three components, $[\hat{p}_{i},\hat{c}_{i},\Delta a^{s}_{i}]$, to encourage the policy to smoothly move toward the target goal pose $p_{g}$ without collision. Here $\hat{p}_{i}$ is the $i$-th virtual robot pose, $\hat{c}_{i}$ is the estimated collision state at $i$-th virtual robot pose $\hat{p}_{i}$ (where zero indicates no collision), and $\Delta a^{s}_{i}$ indicates the action difference, $a^{s}_{i+1}-a^{s}_{i}$. Accordingly, we define $s^{ref}$ as [$p_{g}$, 0.0, 0.0].
@@
The states $\{s^{s}_{i}\}_{i=0\ldots N-1}$ are calculated by computing rollouts through a differentiable forward dynamic model $f$ (in this case, the unicycle model). The forward model considers the current observation $O_{c}$ and generated actions $\{a^{s}_{i}\}_{i=0\ldots N-1}$ from the short-horizon MBRA model, $\pi^{s}$ [29]:
$\{s^{s}_{i}\}_{i=0\ldots N-1}=f(O_{c},\{a^{s}_{i}\}_{i=0\ldots N-1}),$ (2)
@@
where $O_{c}$ is the current observation. While the states $\{s^{s}_{i}\}_{i=0\ldots N-1}$ are conditioned on actions $\{a^{s}_{i}\}_{i=0\ldots N-1}$ and $f$ is differentiable, we can calculate the gradient of $\pi_{s}$ to minimize $J_{mbl}$ in each training step and learn $\pi_{s}$ by repetitively update the parameters of $\pi_{s}$ similar to other machine learning approaches. We do not modify the forward dynamics model $f$ during this training $\pi_{s}$.
@@
Note that while we are still training $\pi^{s}$ on the suboptimal dataset in addition to $\mathcal{D}^{*}$, we are not directly imitating actions. By relying on the forward model $f$ and a reasonable distant target pose $p_{g}$, we can mitigate the effects of both the suboptimal action labels as well as noisy tracking information. Therefore, the MBRA model can leverage the visually and behaviorally diverse dataset, despite the low-quality actions. We discuss implementation details in Section 4.
3.2 Learning a Long-Horizon Navigation Policy, LogoNav
To train our long-horizon navigation policy, we first re-annotate the crowd-sourced dataset $\mathcal{D}_{n}$ with our learned $\pi^{s}$. This gives us a clean set of action labels that can be distilled into an end-to-end navigation policy $\pi^{l}$. We want a navigation policy $\pi^{l}$ to predict actions as:
$\{a^{l}_{i}\}_{i=0\ldots N-1}=\pi^{l}(O_{c},p_{g}),$ (3)
where $O_{c}$ is the current observation and $p_{g}$ is the 2D relative goal pose from the robot coordinate. Notably, $p_{g}$ is at least 10 times further than the usual goal pose for the short-horizon relabeling model, on the order of 50 meters, compared to the previous 3 meters. We train this policy using imitation learning on the re-annotated action commands $\{a^{s}_{i}\}_{i=0\ldots N-1}$ from the short-horizon relabeling model as follows:
$\min\,J_{il}:=\sum_{i=0}^{N-1}(a^{s}_{i}-a^{l}_{i})^{2}.$ (4)
@@ 3.2 Learning a Long-Horizon Navigation Policy, LogoNav
By imitating the cleaned action commands linking $O_{c}$ and $O_{g}$, our long-horizon policy, LogoNav , can learn navigational affordances, such as staying on paths, avoiding collisions, and not disturbing pedestrians, which is representative of the “good” navigation behavior modeled by the MBRA model. Note that we co-train on the relabeled $\mathcal{D}_{n}$ as well as the high-quality dataset $\mathcal{D}^{*}$. We freeze $\pi^{s}$ while training $\pi^{l}$.
4 Implementation
@@
4.1 Passive Dataset
We evaluate our approach with two different passive datasets, a crowd-sourced robotic dataset, FrodoBots-2k, and an in-the-wild YouTube video dataset described in [28]. We focus on results using FrodoBots-2k to demonstrate the effectiveness of our proposed approach and additionally evaluate its capabilities on the YouTube video dataset.
Crowd-sourced robotic dataset: The FrodoBots-2k dataset [30] includes 2000 hours of data from over 10 cities and was collected as part of FrodoBots AI, where users explore locations worldwide by teleoperating robots to reach target positions. The FrodoBots-2k dataset is significantly larger than other publicly available datasets for vision-based navigation tasks. As shown in Table 1, the full version of the FrodoBots-2k is more than 25 times larger than other datasets and includes a diverse set of real robot trajectories teleoperated by humans.
Table 1: Survey of public datasets for learning vision-based navigation policies in real-world.
@@ 4.1 Passive Dataset
KITTI odom [31] teleop 0.7 RGB, 3D LiDAR, GPS
NCLT [32] teleop 34.9 RGB, 3D LiDAR, odom, GPS, IMU
GO Stanford [20, 12] teleop 10.3 RGBs, odom
FLOBOT [33] auto 0.46 RGBD, 3D and 2D LiDAR, odom, IMU.
RECON [14] auto 25.0 stereo RGBD, 2D LiDAR, GPS, IMU
JRDB [34] teleop 1.1 stereo RGBD, 3D and 2D LiDAR, IMU
SCAND [35] teleop 8.7 RGBD, 3D LiDAR, odom
TartanDrive [36] teleop 5.0 RGBD, GPS, IMU
HuRoN [37] teleop 75.0 RGBs, 2D LiDAR, odom, bumper
@@
While the scale and diversity of this dataset are enticing, the inexpensive hardware setup of the robots and crowd-sourcing approach result in significant noise. Since the sensor measurements cannot be reliably used to estimate robot poses, policies trained on the raw actions have poor performance. The main factors of the noisy action labels are 1) robot inconsistencies and corresponding user adjustments, 2) low-cost GPS and IMU, 3) inevitable wheel slips during turning, 4) robot vibration during turning, and 5) system time delay. The details of the robot system and the noisy action labeling are shown in the appendix.
@@
In-the-wild YouTube videos: We also evaluate the ability of MBRA to enable the use of non-robot data. We reannotate 100 hours of action-free in-the-wild YouTube videos, listed in [28], and train a version of LogoNav with the generated actions. These videos include inside and outside walking tours from 32 different countries across varying weather conditions, time of day, and environment types (urban, rural, etc.).
@@
In addition to the passive data, we use the public expert datasets RECON [14], GO Stanford [20, 12], CoryHall [38], TartanDrive [36], HuRoN [37], Seattle [39], and SCAND [35] with accurate action labels. The weighting of each dataset is the same as the original GNM dataset mixture.
4.2 Pre-Processing and Filtering

### v3

@@ Abstract
Developing broadly generalizable visual navigation policies for robots is a significant challenge, primarily constrained by the availability of large-scale, diverse training data. While curated datasets collected by researchers offer high quality, their limited size restricts policy generalization. To overcome this, we explore leveraging abundant, passively collected data sources, including large volumes of crowd-sourced teleoperation data and unlabeled YouTube videos, despite their potential for lower quality or missing action labels. We propose Model-Based ReAnnotation (MBRA), a framework that utilizes a learned short-horizon, model-based expert model to relabel or generate high-quality actions for these passive datasets. This relabeled data is then distilled into LogoNav, a long-horizon navigation policy conditioned on visual goals or GPS waypoints. We demonstrate that LogoNav, trained using MBRA-processed data, achieves state-of-the-art performance, enabling robust navigation over distances exceeding 300 meters in previously unseen indoor and outdoor environments. Our extensive real-world evaluations, conducted across a fleet of robots (including quadrupeds) in six cities on three continents, validate the policy’s ability to generalize and navigate effectively even amidst pedestrians in crowded settings. We present videos showcasing the performance and release our checkpoints and training code on our project page111model-base-reannotation.github.io.
I Introduction
Machine learning has demonstrated remarkable success across a range of tasks, including natural language processing [vaswani2017attention, brown2020language] and computer vision [radford2021learning, kirillov2023segment, oquab2023dinov2]. A key factor driving these advancements is the availability of large and diverse training datasets. In robotics, lack of data is a major bottleneck: intentional, centralized data-collection efforts are costly, requiring real-world robots and human operators, while Internet-scraped data is rarely directly applicable to the robotics domain [o2023open, khazatsky2024droid].
In this paper, we study the problem of developing an end-to-end robot navigation policy capable of generalizing to a wide range of outdoor and indoor environments and navigating to distant goals hundreds of meters away. Training such an end-to-end policy requires large amounts of diverse data to grant broad coverage over possible environments. Previous navigation works [shah2023gnm] have relied on centrally collected datasets generated by robotics researchers. While these datasets tend to be high quality, the sum total of these datasets is on the order of dozens of hours [shah2023vint], limiting the breadth of generalization that can be achieved from this high-quality data alone.
@@ I Introduction
Figure -1: Overview of MBRA. We propose a two-step process: In the first stage, we train a short-horizon reannotation policy with a robust MBL approach on the noisy dataset, which can be used for short-horizon image-conditioned navigation and which we leverage to relabel the noisy dataset with improved action labels. In step 2, we train a long-horizon navigation policy with the generated action labels.
II Related work
Vision-based robot navigation has been widely explored to navigate toward goal positions given visual observations from a monocular camera. [savinov2018semi, pathak2018zero, hirose2019deep] train short-horizon policies to generate actions with access to a single goal observation. These short-horizon policies often utilize topological memory to extend the range of navigation [meng2020largescaletopo]. Some works[shah2021rapid] use exploration with a topological memory to seek out a distant image goal, while others[shah2022viking, stachowicz2023fastrlap] use a GPS signal for localization and navigate toward a goal provided as a 2D position in cartesian coordinates. Goal images and poses require prior access to the target environment and knowledge of the environment’s geometry. Various learning methodologies such as imitation learning (IL) [savinov2018semi, shah2023gnm, shah2023vint], reinforcement learning (RL) [stachowicz2023fastrlap, hiroseselfi, stachowicz2024lifelong], and model-based learning (MBL) [hirose2019deep, hirose2023exaug] have been explored for training goal-conditioned vision-based policies on publicly available robot datasets.
These methods require a sequence of image observations and corresponding actions parsed from accurate wheel odometry [hirose2019deep, hirose2018gonet], GPS [shah2021rapid], and other reliable sensors. These datasets are collected via intentional, centralized teleoperation efforts with the downstream goal of training a navigation policy and, therefore, contain goal-directed trajectories. Collecting data of this sort at a global scale would require a massive unified effort that would be costly and time-consuming.
Robot learning with passive data. Visual SLAM [liu2024citywalker] and inverse dynamics models (VPT) [baker2022videopretrainingvptlearning] can be used to estimate trajectories for first-person videos, allowing us to train policies that use these trajectories as approximations of robot actions from action-free and non-robot data. While visual SLAM and its successors [orbslam3, mast3rslam, droidslam, dpvo] offer good local trajectory estimation, its accuracy relies on having consistent, good visual features in the image view.
Prior methods have also sought to leverage suboptimal data, including data without actions, by training a separate model to infer the action given the current and next state [baker2022videopretrainingvptlearning]. We find that this approach performs worse than MBRA on the highly suboptimal datasets we consider, both because the action prediction must leverage other datasets that contain distributional shift. Perhaps even more importantly, MBRA generates synthetic actions that are optimized to reach future states in a trajectory while satisfying basic navigational objectives via a forward model, rather than greedily maximizing the probability of the action given a pair of adjacent states, leading to smoother and more reasonable behavior.
Related to inverse model methods such as VPT, multi-step goal-conditioned policies [shah2023vint, sridhar2024nomad] (GCPs) train a model to predict the action given the current state and a (more distant) future state. While these approaches can also take non-greedy actions, unlike MBRA, they do not optimize for actions that satisfy navigational objectives. They also still suffer from distributional shift when the action labels in the target domain are unavailable, or else must use noisy action labels if they are present, both of which degrade performance (see Fig. II[b]). In our experiments, we find that MBRA significantly outperforms such methods when using large datasets with low-quality actions.
Figure 0: The aim of MBRA is to relabel low-quality data with actions that are better than the actions in the dataset, in the sense that they more effectively link states over short-horizon trajectory snippets. Compared to methods that use one- or multi- step inverse models (e.g., VPT, GCP) or the original noisy actions, training on actions from MBRA leads to significantly more effective policies.
@@ II Related work
III Learning Short-Horizon Relabeling Policies with Model-based Learning
@@
In this paper, we focus on learning a long-horizon navigation policy from a highly diverse but suboptimal dataset $\mathcal{D}_{n}$. In particular, we wish to learn high-quality navigation from a crowd-sourced dataset; this requires us to train a relabeler that can predict actions that are better than those found in the original dataset. We assume access to a smaller clean dataset $\mathcal{D}^{*}$ that contains high-quality behavior, such that $\lvert\mathcal{D}^{*}\rvert<\!\!<\lvert\mathcal{D}_{n}\rvert$.
While observations in $\mathcal{D}_{n}$ might represent high state coverage, the actions are low-quality: both because of inaccuracy due to state estimation errors and heterogeneity of uncurated human operators with varying skill levels. Our key insight is that a short-horizon model-based expert trained from these noisy datasets can be used to relabel entire trajectories. This leads to a high-quality training dataset for an end-to-end navigation policy that can imitate these clean actions (Fig. -1).
III-A Learning a Short-Horizon Relabeling Model, MBRA
To train an accurate MBRA model $\{a^{s}_{i}\}_{i=0\ldots N-1}=\pi^{s}(O_{c};O_{g})$ to infer the optimal actions occurring between the current observation $O_{c}$ and the goal observation $O_{g}$ for re-annotation, we propose the robust MBL to leverage the entire suboptimal dataset $\mathcal{D}_{n}$ as well as $\mathcal{D}^{*}$ in training.
Overview of learning MBRA model. As illustrated in Fig. II[b], single-step VPT and multi-step GCPs struggle to learn from noisy action labels during training. To address this, we use Model-based Learning (MBL). MBL prioritizes the final goal state, but rather than mimicking the potentially noisy actions in the dataset, it utilizes a forward model to generate synthetic actions. Crucially, these actions are optimized both to reach future states in the trajectory and to satisfy basic navigational constraints (as shown in Fig. II[c]). This approach leads to smoother and more reasonable behavior, allowing MBL to leverage the noisy FrodoBots-2k more effectively and resulting in better labels for the entire FrodoBots-2k dataset. This allows us to improve the action label quality, while also preserving teleoperators’ intent, such as avoiding collisions with pedestrians, staying on paths, etc.
Learning architecture based on MBL. We design the following model-based objective $J_{mbl}$ for learning $\pi^{s}$ to reach target state $s^{ref}$ with keeping basic navigation constraints such as avoiding collision. Note that we only give the further target state $s^{ref}$ instead of giving the individual adjacent target states such as $s^{ref}_{i}$ in each $i$-th step not to be sensitive for low-quality states in the dataset as shown in Fig. II[b].
@@ III-A Learning a Short-Horizon Relabeling Model, MBRA
where $\{s^{s}_{i}\}_{i=0\ldots N-1}$ are the estimated states at each step. The states $\{s^{s}_{i}\}_{i=0\ldots N-1}$ are calculated by computing rollouts through a differentiable forward dynamic model $f$. The forward model considers the current observation $O_{c}$ and generated actions $\{a^{s}_{i}\}_{i=0\ldots N-1}$ from the short-horizon MBRA model, $\pi^{s}$ [hirose2022depth360].
$\{s^{s}_{i}\}_{i=0\ldots N-1}=f(O_{c},\{a^{s}_{i}\}_{i=0\ldots N-1}).$ (2)
@@
While the states $\{s^{s}_{i}\}_{i=0\ldots N-1}$ are conditioned on actions $\{a^{s}_{i}\}_{i=0\ldots N-1}$ and $f$ is differentiable, we can calculate the gradient of $\pi_{s}$ to minimize $J_{mbl}$ in each training step and learn $\pi_{s}$ by repetitively update the parameters of $\pi_{s}$ similar to other machine learning approaches. We do not modify the forward dynamics model $f$ during this training $\pi_{s}$. Detail implementation is shown in the Sec.V-E.
@@
Note that while we are still training $\pi^{s}$ on the suboptimal dataset in addition to $\mathcal{D}^{*}$, we are not directly imitating actions. By relying on the forward model $f$ and a reasonable distant target state $s^{ref}$, we can mitigate the effects of both the suboptimal action labels as well as noisy tracking information (Fig. II[c]). Therefore, the MBRA model can leverage the visually and behaviorally diverse dataset, despite the low-quality actions.
III-B Learning a Long-Horizon Navigation Policy, LogoNav
@@
To train our long-horizon navigation policy, we first re-annotate the crowd-sourced dataset $\mathcal{D}_{n}$ with our learned $\pi^{s}$. This gives us a clean set of action labels that can be distilled into an end-to-end navigation policy $\pi^{l}$. We want a navigation policy $\pi^{l}$ to predict actions as $\{a^{l}_{i}\}_{i=0\ldots N-1}=\pi^{l}(O_{c},p_{g})$ where $O_{c}$ is the current observation and $p_{g}$ is the 2D relative goal pose from the robot coordinate. Notably, $p_{g}$ is at least 10 times further than the usual goal pose for the short-horizon relabeling model, on the order of 50 meters, compared to the previous 3 meters. We train this policy using imitation learning on the re-annotated action commands $\{a^{s}_{i}\}_{i=0\ldots N-1}$ from the short-horizon relabeling model such as $\min\,J_{il}:=\sum_{i=0}^{N-1}(a^{s}_{i}-a^{l}_{i})^{2}$. By imitating the cleaned action commands linking $O_{c}$ and $O_{g}$, our long-horizon policy, LogoNav , can learn navigational affordances, such as staying on paths, avoiding collisions, and not disturbing pedestrians, which is representative of the “good” navigation behavior modeled by the MBRA model. Note that we co-train on the relabeled $\mathcal{D}_{n}$ as well as the high-quality dataset $\mathcal{D}^{*}$. We freeze $\pi^{s}$ while training $\pi^{l}$.
IV Implementation
@@
IV-A Passive Dataset
We evaluate our approach with two different passive datasets, a crowd-sourced robotic dataset, FrodoBots-2k, and an in-the-wild YouTube video dataset described in [hiroselelan]. We focus on results using FrodoBots-2k to demonstrate the effectiveness of our proposed approach and additionally evaluate its capabilities on the YouTube video dataset.
TABLE I: Survey of public datasets for learning vision-based navigation policies in real-world.
@@ IV-A Passive Dataset
KITTI odom [Geiger2012CVPR] teleop 0.7 RGB, 3D LiDAR, GPS
NCLT [carlevaris2016university] teleop 34.9 RGB, 3D LiDAR, odom, GPS, IMU
GO Stanford [hirose2018gonet, hirose2019deep] teleop 10.3 RGBs, odom
FLOBOT [yan2020robot] auto 0.46 RGBD, 3D and 2D LiDAR, odom, IMU.
RECON [shah2021rapid] auto 25.0 stereo RGBD, 2D LiDAR, GPS, IMU
JRDB [martin2021jrdb] teleop 1.1 stereo RGBD, 3D and 2D LiDAR, IMU
SCAND [karnan2022socially] teleop 8.7 RGBD, 3D LiDAR, odom
TartanDrive [triest2022tartandrive] teleop 5.0 RGBD, GPS, IMU
HuRoN [hirose2023sacson] teleop 75.0 RGBs, 2D LiDAR, odom, bumper
@@
Crowd-sourced robotic dataset: The FrodoBots-2k dataset [frodobots_2K] includes 2000 hours of data from over 10 cities and was collected as part of FrodoBots AI, where users explore locations worldwide by teleoperating robots to reach target positions. The FrodoBots-2k dataset is significantly larger than other publicly available datasets for vision-based navigation tasks. As shown in Table I, the full version of the FrodoBots-2k is more than 25 times larger than other datasets and includes a diverse set of real robot trajectories teleoperated by humans.
@@
While the scale and diversity of this dataset are enticing, the inexpensive hardware setup of the robots and crowd-sourcing approach result in significant noise. Since sensor measurements cannot be reliably used to estimate robot poses, policies trained on raw actions have poor performance. The main factors of noisy action labels are 1) robot inconsistencies and corresponding user adjustments, 2) low-cost GPS and IMU, 3) inevitable wheel slips during turning, 4) robot vibration during turning, and 5) system delay. Details of the robot system are shown in [erz_sdk] and Sec. V-A.
@@
In-the-wild YouTube videos: We also evaluate the ability of MBRA to enable the use of non-robot data. We reannotate 100 hours of action-free in-the-wild YouTube videos, listed in [hiroselelan], and train a version of LogoNav with the generated actions. These videos include inside and outside walking tours from 32 different countries across varying weather conditions, time, and environment types (urban, rural, etc.).
@@
In addition to the passive data, we use the public expert datasets RECON [shah2021rapid], GO Stanford [hirose2018gonet, hirose2019deep], CoryHall [kahn2018self], TartanDrive [triest2022tartandrive], HuRoN [hirose2023sacson], Seattle [shaban2022semantic], and SCAND [karnan2022socially] with accurate action labels. The weighting of each dataset is the same as the original GNM dataset mixture.
IV-B Pre-Processing and Filtering